OpenStudy (anonymous):

If a rectangle has a perimeter of 70, a length of x and a width of x - 9, find the value of the length of the rectangle. 4 31 13 22

OpenStudy (anonymous):

22

OpenStudy (anonymous):

perimeter =2(length+width)

OpenStudy (jh3power):

it's 2x+2y=70 so x+y=35 and y=x-9 so it makes 2x-9=35 2x=44 x=22